b. Calculate the coefficient of price elasticity of demand from the given data for point A to B.

Point Quantity Price

A 4 7

B 3 8



1
Expert's answer
2021-04-20T10:43:39-0400

The formula for the coefficient of PED is:


PED="\frac{ (percentage)change in quantity demanded}{(percentage)change in price}"


%change in quantity demanded =3-4=-1

"=\frac{-1}{3}×" 100%=-33.3%


%Change in price =8-7=1

"=\frac{1}{8}×" 100%=12.5%


"\therefore" Coefficient of PED="\frac{-33.3}{12.55}"


="-2.7"

="2.7"